# Hierarchical Large Scale Multirobot Path (Re)Planning

### v1

@@ Abstract
We consider a large-scale multi-robot path planning problem in a cluttered environment. Our approach achieves real-time replanning by dividing the workspace into cells and utilizing a hierarchical planner. Specifically, multi-commodity flow-based high-level planners route robots through the cells to reduce congestion, while an anytime low-level planner computes collision-free paths for robots within each cell in parallel. Despite resulting in longer paths compared to the baseline multi-agent pathfinding algorithm, our method produces a solution with significant improvement in computation time. Specifically, we show empirical results of a $500$-times speedup in computation time compared to the baseline multi-agent pathfinding approach on the environments we study. We account for the robot’s embodiment and support non-stop execution when replanning continuously. We demonstrate the real-time performance of our algorithm with up to $142$ robots in simulation, and a representative $32$ physical Crazyflie nano-quadrotor experiment.
@@ I Introduction
Large fleets of robots, such as those used in warehouse operations [1], disaster response [2], and delivery [3], demand coordination solutions that adjust in real time to changing goals. In this work, we present a real-time lifelong hierarchical method for navigating a large team of robots to independent goals in a large, cluttered environment that guarantees collision avoidance. By lifelong, we mean robots can enter and exit the space, and can receive another goal at any time, as they would in a warehouse or delivery problem. Our approach partitions the space into disjoint cells, allowing planning algorithms to run concurrently in parallel within each cell. A high-level planner routes robots through the partition, while a low-level anytime multi-agent pathfinding (MAPF) algorithm navigates robots to local goals within each cell in parallel. The real-time property holds as long as there are not too many cells or robots in a workspace; the limits for real-time operation are empirical and problem-specific, however, we demonstrate real-time performance for 142 robots in simulation with a 25-cell partition.
@@
We are particularly interested in unmanned aerial vehicles (UAVs) operating in 3D space, such as in city-scale on-demand UAV package delivery, however, our approach applies to robots operating in 2D as well. We present two approaches for high-level planning depending on the problem’s requirements: 1) an egocentric greedy approach that always operates in real-time and 2) a novel high-level planner that routes robots through the partition using multi-commodity flow (MCF) [4]. There are tradeoffs between these two approaches. The egocentric greedy planner operates in real-time regardless of the number of cells; however, it has no mechanism for distributing robots, thus it can result in congestion and longer low-level planning times within some cells. On the other hand, the MCF-based approach eases cell congestion by regulating the flow of robots into each cell while ensuring bounded-suboptimal inter-cell routing; thus, it can be useful in environments such as urban UAV package delivery, where different types of cells (e.g., residential vs. highway) may have different limits on the influx of robots. The MCF-based planners can operate in real-time under certain conditions, thus allowing for lifelong replanning while reducing congestion, which leads to faster, real-time low-level planning within each cell.
The low-level planner ensures collision avoidance while respecting the robots’ geometric shape. A cell-crossing protocol allows robots to transition between cells without stopping in midair. Combined with the MCF-based planner, this allows real-time computation and safe, non-stop execution of multi-robot plans. The contributions of this work are:
@@ II Related Work
Centralized approaches to multi-robot planning [5, 6] face substantial computational challenges due to the theoretical hardness of the problem [7], prohibiting real-time replanning and scaling to many robot systems. Decentralized approaches, on the other hand, can result in deadlocks, livelocks, congestion, collision, and reduced efficiency. RLSS [8] uses fast single-robot planners but relies on additional optimization to provide collision-free trajectories. In cluttered environments, a buffered Voronoi cell based algorithm [9] can result in deadlocks, and an algorithm using relative safe flight corridor [10] leads to collisions. In the present work, we aim to address these problems and facilitate real-time MAPF at the discrete planning phase.
Search-based MAPF solvers generate high-quality, collision-free discrete paths but the complexity for optimal solutions scales exponentially with the number of agents [7]. Bounded-suboptimal algorithms have been proposed to overcome this complexity, but their poor scalability still prevents their application to real-time coordination for large teams. To address this, partition-based MAPF [11, 12] divides the workspace into smaller regions, reducing the agent number within each cell. However, since the high-level planners are either single-agent based [11] or solving a multi-commodity flow problem constrained on single-robot shortest paths [12], they lead to congestion and hard MAPF instances. Instead, we propose a novel inter-cell routing algorithm to distribute robots while maintaining bounded-suboptimality. Furthermore, the cell-crossing method in [12] is unsuitable for aerial vehicles due to energy consumption while waiting in place for the cell-crossing channel to clear. Our cell-crossing protocol addresses this drawback.
@@
III Problem Formulation
Consider a time-varying number of homogeneous non-point robots operating in workspace $\mathcal{W}$, which is partitioned into a union of non-overlapping convex polytopic cells. Robots must reach specified individual goal positions, which change over time, while avoiding collisions with robots and obstacles and obeying maximum cell influx limits $\boldsymbol{\theta}$ (influx refers to the number of robots entering a workspace cell). A motivating scenario is a multi-UAV package delivery system where the number of UAVs entering some types of airspace must be limited, and UAVs exit or enter the workspace to charge or redeploy.
The problem above requires solving a flow problem that obeys cell influx limits while handling new goal positions and a varying number of robots. While that is the general problem we aim to solve, the present work addresses a critical subproblem: the underlying planner that is repeatedly called to safely and efficiently route the robots through $\mathcal{W}$.
Consider $N$ (fixed) homogeneous non-point robots operating in a workspace $\mathcal{W}$ that is partitioned into a union of non-overlapping convex polytopic cells. $\mathcal{W}$ contains obstacles represented as unions of convex polytopes $\mathcal{O}_{1},\cdots,\mathcal{O}_{N_{obs}}$. Let $\mathcal{R}_{\mathcal{E}}(\mathbf{p})$ be the convex set of points representing a robot at position $\mathbf{p}$, i.e., a robot-environment collision model. The free space is represented as $\mathcal{F}=\mathcal{W}\backslash(\bigcup_{h}\mathcal{O}_{h})\ominus\mathcal{R}_{\mathcal{E}}(\mathbf{0})$, where $\ominus$ denotes Minkowski difference. For each robot $r^{i}$ (superscript $i$ represents robot index) with initial position $\mathbf{s}^{i}\in\mathbb{R}^{3}$, find paths for all robots to their goal positions such that there are no collisions (e.g., between robots or between robots and $\bigcup_{h}\mathcal{O}_{h}$) and the total number of robots that enter each cell is less than its user-defined influx $\theta_{m}$ (influx limits can vary by cell).
IV Preliminaries
IV-A Multi-agent pathfinding (MAPF)
@@
Consider an undirected graph $\mathcal{G}\!\!=\!\!\left(V,E\right)$, and $N$ agents. Each agent has a start $v^{i}_{s}\!\!\in\!\!V$ and goal $v^{i}_{g}\!\!\in\!\!V$ vertex. At each time step $k$, an agent can either move to a neighbor vertex $\left(u^{i}_{k},u^{i}_{k+1}\right)\!\!\in\!\!E$ or stay at its current vertex $u^{i}_{k+1}\!\!=\!\!u^{i}_{k}$, where $u^{i}_{k}\!\!\in\!\!V$ is the $k$-th vertex in $i$-th agent’s path. To respect vertex conflict constraints, no two agents can occupy the same vertex simultaneously, i.e., $\forall k,i\!\!\neq\!\!j\!\!:\!\!u^{i}_{k}\neq u^{j}_{k}$. To respect edge conflict constraints, no two agents can traverse the same edge in the opposite direction concurrently, i.e., $\forall k,i\!\!\neq\!\!j\!\!:\!\!u^{i}_{k}\neq u^{j}_{k+1}\!\vee\!u^{i}_{k+1}\!\!\neq\!\!u^{j}_{k}$. The objective is to find conflict-free paths $\mathcal{P}^{i}\!\!=\!\!\left[u^{i}_{0},\cdots,u^{i}_{T-1}\right]$, where $u^{i}_{0}\!\!=\!\!v^{i}_{s}$ and $u^{i}_{T-1}\!\!=\!\!v^{i}_{g}$ for all agents, and minimize cost, e.g., the sum over the time steps required to reach the goals of all agents or the makespan $T$.
IV-B Conflict Annotation and MAPF with General Conflicts
@@
Many works address MAPF for embodied agents [13, 14, 15]. We adopt multi-agent pathfinding with generalized conflicts (MAPF/C), due to its flexibility in incorporating different geometric shapes [5]. Conflict annotation identifies the extended conflict set of vertices and edges for each vertex and edge with respect to the inter-robot collision model. To account for the downwash effect between robots [16], we use a robot-robot collision model $\mathcal{R}_{\mathcal{R}}(\mathbf{p})$ that is distinct from $\mathcal{R}_{\mathcal{E}}(\mathbf{p})$. We follow the MAPF/C function definitions [5]:
@@
where $\operatorname{pos}(u)\in\mathbb{R}^{3}$ returns the position for vertex $u$. $\mathcal{R}_{\mathcal{R}}^{*}(e)$ is the set of points swept by the robot when traversing edge $e$. This is called the swept collision model.
@@ V Geometric Partitioning
Our framework has three components: geometric partitioning, high-level planner, and low-level planner, as depicted in Fig. 2. Geometric partitioning divides the workspace into disjoint convex cells, where the plan can be computed in parallel. A centralized high-level planner regulates the congestion for each cell while guaranteeing robots’ inter-cell routing quality. An anytime low-level planner plans collision-free paths for robots within each cell. Initial planning includes pre-computation (the geometric partitioning), and replanning only involves high- and low-level planning. Once the initial plan is established, our algorithm runs in real-time for replanning.
@@ V-A Roadmap Generation
A roadmap is an undirected graph $\mathcal{G}\!\!=\!\!(V,E)$ embedded into a Euclidean space, where each vertex $v\!\!\in\!\!V$ corresponds to a position in $\mathcal{F}$ and each edge $(u,v)\!\!\in\!\!E$ denotes a path in $\mathcal{F}$ connecting $u$ and $v$. We additionally require the existence of vertices $v^{i}_{g}$ and $v^{i}_{s}$, corresponding to the goal and start positions, $\mathbf{g}^{i}$ and $\mathbf{s}^{i}$, respectfully. The roadmap should satisfy three properties: 1) connectivity-preserving, i.e., if a path between two points in $\mathcal{F}$ exists, there should be a path in the roadmap as well; 2) optimality-preserving, i.e., the shortest path between two points in $\mathcal{F}$ can be well approximated by a path in the roadmap; and 3) sparse, i.e., have a small number of vertices and edges. In our experiments, we use a 6-connected grid graph as roadmap, however, it can be generated by other methods, such as SPARS [17].
V-B Graph Partitioning and Spatial Linear Separation
Our method partitions the workspace into disjoint convex cells and within each cell solves a MAPF instance. Spatial decomposition has two benefits: 1) fewer robots for each smaller sub-graph, and 2) the decomposed MAPF instances can be solved in parallel. The initial partitioning can adopt any approach. We propose a geometric method that generates $Q$ convex polytopes. First, we use graph partitioning (KaHyPar [18]) to group the roadmap into $Q$ balanced sub-graphs $\mathcal{G}_{m}\!\!=\!\!\left({V_{m},E_{m}}\right)$, for $m\!\!=\!\!1,\ldots,Q$. Balanced means the vertex number difference between sub-graphs is bounded, leading to cells of similar volume and an even spread of robots.
Despite the sub-graphs now being balanced, we further enforce each cell, containing a sub-graph, to be a convex polytope. To do so, we use soft-margin support vector machines (SVM) [19] to compute the separating hyperplane $H_{ml}$ between sets of vertices of $\mathcal{G}_{m}$ and $\mathcal{G}_{l}$, and reassign misclassified vertices. The resulting set of separating hyperplanes form a cell. Convexity prevents robots from penetrating into neighboring regions while traversing within a cell, thus isolating cell planning.
@@ V-C Local Goal Generation
For navigating out of a cell, we generate candidate goal states on the faces between adjacent cells. For each cell $P_{m}$, we uniformly sample random local goals on the hyperplane $H_{ml}$ and add them as shared vertices to both $\mathcal{G}_{m}$ and $\mathcal{G}_{l}$. Despite being shared vertices, local goals generated by partition $P_{m}$ have in-edges from $P_{m}$ and out-edges to $P_{l}$ to avoid collision during cell transit. Thus, this part of the graph is directed. To enable parallel computation, there must be no communication between cells. Thus, the cell roadmap $\mathcal{G}_{m}$ is modified such that the planned paths are collision-free when crossing cells without information exchange between cells. Thus, the following properties should be satisfied:
P1: Robots avoid collision when stationary at vertices (local goals and non-local-goal vertices) of different cells (generalized vertex-vertex conflict across cells), i.e., $\forall i,j,m\!\!\neq\!\!l\!:\!v^{i}_{m}\!\!\notin\!\!\operatorname{conVV}(v^{j}_{l})$, where the superscript in the $v^{i}_{m}$ refers to the vertex index and the subscript refers to cell index.
P2: Robots avoid collision when traversing edges within different cells (generalized edge-edge conflict across cells), i.e., $\forall i,k,m\!\!\neq\!\!l:e^{i}_{m}\!\!:=\!\!(v^{i}_{m},v^{j}_{m})\notin\operatorname{conEE}(e^{k}_{l})$, here the superscript in $e^{i}_{m}$ refers to the edge index.
@@
We depict violations of P1-P3 in Fig. 3b-3d. To prevent conflicts between local goals and non-local-goal vertices across cells, we buffer the separating hyperplane by the inter-robot collision configuration, $\mathcal{C}_{col}$ (c.f. Fig. 3a), which is computed as $\mathcal{C}_{col}(\mathbf{p})=\mathcal{R}_{\mathcal{R}}(\mathbf{p})\oplus\mathcal{R}_{\mathcal{R}}(\mathbf{0})$, where $\oplus$ is the Minkowski sum. The buffering is achieved by modifying the offset $\mathcal{H}_{a}^{{}^{\prime}}=\mathcal{H}_{a}+\mathrm{max}_{\mathbf{y}\in\mathcal{C}_{col}(\mathbf{0})}\mathcal{H}_{\mathbf{n}}\!\cdot\!\mathbf{y}$ of the hyperplane, where $\mathcal{H}_{a}$ and $\mathcal{H}_{\mathbf{n}}$ are the offset and the normal vector of the hyperplane. Given the buffered hyperplanes, all non-local-goal vertices within the buffered region are removed, guaranteeing collision avoidance between local goals and non-local-goal vertices. The local goals sampled on the hyperplane are confined within the blue region in Fig. 3a to avoid vertex-vertex conflicts between local goals of different separating hyperplanes. The buffering satisfies P1 between non-local-goal and local-goal vertices across cells and P1, P2, and P3 between non-local-goal vertices across cells. To satisfy P1 between local goals across cells, we uniformly randomly sample points on the hyperplane and reject those that violate P1. For P2 and P3 between non-local-goal and local goal vertices, we add the connection between the sampled local goal to its neighbor non-local-goal vertices within a radius on both sides of the hyperplane. The local goal is removed if any connected edge violates P2 or P3 (see Fig. 3c, 3d). Otherwise, the edges are added to the cell roadmaps. Figure 4 depicts an exemplar geometric partition.

### v2

@@ Abstract
We consider a large-scale multi-robot path planning problem in a cluttered environment. Our approach achieves real-time replanning by dividing the workspace into cells and utilizing a hierarchical planner. Specifically, we propose novel multi-commodity flow-based high-level planners that route robots through cells with reduced congestion, along with an anytime low-level planner that computes collision-free paths for robots within each cell in parallel. A highlight of our method is a significant improvement in computation time. Specifically, we show empirical results of a 500-times speedup in computation time compared to the baseline multi-agent pathfinding approach on the environments we study. We account for the robot’s embodiment and support non-stop execution with continuous replanning. We demonstrate the real-time performance of our algorithm with up to 142 robots in simulation, and a representative 32 physical Crazyflie nano-quadrotor experiment.
@@ I Introduction
Large fleets of robots, such as those used in warehouse operations [1], disaster response [2], and delivery [3], demand coordination solutions that adjust in real time to changing goals. In this work, we present a real-time lifelong hierarchical method for navigating a large team of robots to independent goals in a large, cluttered environment that guarantees collision avoidance. By lifelong, we mean robots can enter and exit the space, and can receive another goal at any time, as they would in a warehouse or delivery problem. Our approach partitions the space into disjoint cells, allowing planning algorithms to run concurrently in parallel within each cell. A novel high-level planner routes robots through the partition, while a low-level anytime multi-agent pathfinding (MAPF) algorithm navigates robots to local goals within each cell in parallel. The real-time property holds as long as there are not too many cells or robots in a workspace; the limits for real-time operation are empirical and problem-specific, however, we demonstrate real-time performance for 142 robots in simulation with a 25-cell partition.
@@
We are primarily interested in unmanned aerial vehicles (UAVs) operating in 3D space, such as in city-scale on-demand UAV package delivery; however, our approach applies to robots operating in 2D as well. We present two approaches for high-level planning depending on the problem’s requirements: 1) an egocentric greedy approach that always operates in real-time and 2) a novel high-level planner that routes robots through the partition using multi-commodity flow (MCF) [4]. There are tradeoffs between these two approaches. The egocentric greedy planner operates in real-time regardless of the number of cells; however, it has no mechanism for distributing robots, thus it can result in congestion and longer low-level planning times within some cells. On the other hand, the MCF-based approach eases cell congestion by regulating the flow of robots into each cell while ensuring bounded-suboptimal inter-cell routing; thus, it can be useful in environments such as urban UAV package delivery, where different types of cells (e.g., residential vs. highway) may have different limits on the influx of robots. The MCF-based planners can operate in real-time under certain conditions, thus allowing for lifelong replanning while reducing congestion, which leads to faster, real-time low-level planning within each cell.
The low-level planner prohibits collisions while respecting the robots’ geometric shapes. A novel cell-crossing protocol allows robots to transition between cells without stopping in midair. Combined with the MCF-based planner, this allows real-time computation and safe, non-stop execution of multi-robot plans. The contributions of this work are:
@@ II Related Work
Centralized approaches to multi-robot planning [5, 6] face substantial computational challenges due to their theoretical hardness [7], prohibiting real-time replanning for many-robot systems. Decentralized online approaches, on the other hand, can result in deadlocks, livelocks, congestion, collision, and reduced efficiency. RLSS [8] generates trajectories based on a single-robot planner without a deadlock-free guarantee. Furthermore, control barrier functions [9], or reactive control synthesis methods, are prone to deadlocks. Distributed MPC [10] results in deadlocks when narrow corridors are present. In cluttered environments, a buffered Voronoi cell-based algorithm [11] also suffers from potential deadlocks, and an algorithm using relative safe flight corridor [12] leads to collisions. In the present work, we aim to address these problems and facilitate real-time MAPF at the discrete planning phase.
Search-based MAPF solvers generate deadlock- and collision-free paths, but the complexity of optimal solutions scale exponentially with the number of agents [7]. Bounded-suboptimal algorithms have been proposed to overcome this complexity, but poor scalability still prevents their application to real-time coordination for large teams. To address this, partition-based MAPF [13, 14] divides the workspace into cells, reducing the complexity in each cell. However, inter-cell routing is either single-agent based [13] or the solution of a multi-commodity flow problem constrained on single-robot shortest paths [14] without congestion-awareness, leading to hard instances within regions. Instead, we propose a novel congestion-aware inter-cell routing algorithm to distribute robots while maintaining bounded-suboptimality. Furthermore, the cell-crossing method in [14] is unsuitable for aerial vehicles due to energy expenditure while stationed in hover for the cell-crossing channel to clear. Our novel cell-crossing protocol addresses this drawback.
@@
III Preliminaries
III-A Multi-agent Pathfinding (MAPF) on Euclidean Graph
@@
Consider an undirected graph $\mathcal{G}\!\!=\!\!\left(V,E\right)$ embedded in a Euclidean space, where each vertex $v\!\!\in\!\!V$ corresponds to a position in the free space $\mathcal{F}$ and each edge $(u,v)\!\!\in\!\!E$ denotes a path in $\mathcal{F}$ connecting vertices $u$ and $v$. For $N$ agents, we additionally require the existence of vertices $v^{i}_{g}$ and $v^{i}_{s}$, corresponding to the goal and start positions, $\mathbf{g}^{i}$ and $\mathbf{s}^{i}\!\!\in\!\!\mathbb{R}^{3}$ of robot $r^{i}$ (superscript $i$ represents robot index), respectfully. At each time step $k$, an agent can either move to a neighbor vertex $\left(u^{i}_{k},u^{i}_{k+1}\right)\!\!\in\!\!E$ or stay at its current vertex $u^{i}_{k+1}\!\!=\!\!u^{i}_{k}$, where $u^{i}_{k}\!\!\in\!\!V$ is the occupied vertex in the $i$-th agent’s path at time step $k$. To respect vertex conflict constraints, no two agents can occupy the same vertex simultaneously, i.e., $\forall k,i\!\!\neq\!\!j\!\!:\!\!u^{i}_{k}\neq u^{j}_{k}$. To respect edge conflict constraints, no two agents can traverse the same edge in the opposite direction concurrently, i.e., $\forall k,i\!\!\neq\!\!j\!\!:\!\!u^{i}_{k}\neq u^{j}_{k+1}\!\vee\!u^{i}_{k+1}\!\!\neq\!\!u^{j}_{k}$. The objective is to find conflict-free paths $\mathcal{P}^{i}\!\!=\!\!\left[u^{i}_{0},\cdots,u^{i}_{T-1}\right]$, where $u^{i}_{0}\!\!=\!\!v^{i}_{s}$ and $u^{i}_{T-1}\!\!=\!\!v^{i}_{g}$ for all agents, and minimize cost, e.g., the sum over the time steps required to reach the goals of all agents or the makespan $T$.
III-B MAPF for Embodied Agents and Conflict Annotation
@@
Many works address MAPF for embodied agents [15, 16, 17]. We adopt multi-agent pathfinding with generalized conflicts (MAPF/C), due to its flexibility with different shapes. Generalized conflicts, different from typical MAPF conflicts, include the extra conflicts caused by the robot embodiment [5]. To account for the downwash effect between robots [18], we denote $\mathcal{R}_{\mathcal{R}}(\mathbf{p})$ as the convex set of points representing a robot at position $\mathbf{p}$, i.e., a robot-robot collision model. We follow the conflict annotation in [5] to annotate the graph with generalized conflicts, defined as:
@@
where $\operatorname{pos}(u)\!\!\in\!\!\mathbb{R}^{3}$ returns the position of vertex $u$. $\mathcal{R}_{\mathcal{R}}^{*}(e)$ is the swept collision model representing the set of points swept by the robot when traversing edge $e$.
IV Problem Formulation
Consider a time-varying number of homogeneous non-point robots in workspace $\mathcal{W}$, which is partitioned into a union of disjoint convex polytopic cells. Robots must reach specified individual goal positions, which change over time, while avoiding collisions with robots and obstacles and obeying maximum cell influx limits $\boldsymbol{\theta}$ (influx refers to the number of robots entering a workspace cell). A motivating scenario is a multi-UAV package delivery system, where the number of UAVs entering certain types of airspaces must be limited, and UAVs exit or enter the workspace to charge or redeploy. The problem above requires solving a path replanning problem that obeys cell influx limits while handling new goal positions and a varying number of robots. While that is the general problem we aim to solve, the present work addresses a critical subproblem: the underlying planner that is repeatedly called to safely and efficiently route the robots through $\mathcal{W}$.
Now, consider $N$ (fixed) homogeneous non-point robots operating in the partitioned workspace. $\mathcal{W}$ contains obstacles defined as unions of convex polytopes $\mathcal{O}_{1},\cdots,\mathcal{O}_{N_{obs}}$. We use a robot-environment collision model $\mathcal{R}_{\mathcal{E}}(\mathbf{p})$ that is distinct from $\mathcal{R}_{\mathcal{R}}(\mathbf{p})$. The free space is $\mathcal{F}\!\!=\!\!\mathcal{W}\backslash(\bigcup_{h}\mathcal{O}_{h})\ominus\mathcal{R}_{\mathcal{E}}(\mathbf{0})$, where $\ominus$ is Minkowski difference. For each robot $r^{i}$ at initial position $\mathbf{s}^{i}$, our algorithm finds a path to its goal position $\mathbf{g}^{i}$ such that there are no collisions (e.g., between robots or between robots and $\bigcup_{h}\mathcal{O}_{h}$) and the total number of robots that enter each cell is less than its user-defined influx $\theta_{m}$ (influx limits can vary by cell).
To efficiently address the above problem, our framework has three components: geometric partitioning, high-level planner, and low-level planner, as depicted in Fig. 2. Geometric partitioning divides the workspace into disjoint convex cells, where the plan can be computed in parallel. A centralized high-level planner regulates the congestion for each cell while guaranteeing inter-cell routing quality. An anytime low-level planner plans collision-free paths for robots within each cell. Initial planning includes pre-computation (the geometric partitioning), and replanning only involves high- and low-level planning. Once the initial plan is established, our algorithm runs in real-time for replanning.
@@ V-A Roadmap Generation
A roadmap is an undirected graph, introduced in Sec.III-A, satisfying three properties: 1) connectivity-preserving, i.e., if a path between two points in $\mathcal{F}$ exists, there should be a path in the roadmap as well; 2) optimality-preserving, i.e., the shortest path between two points in $\mathcal{F}$ can be well approximated by a path in the roadmap; and 3) sparse, i.e., have a small number of vertices and edges. In our experiments, we use a 6-connected grid graph, however, it can be generated by other methods, such as SPARS [19].
@@
V-B Graph Partitioning and Spatial Linear Separation
Our method partitions the workspace into disjoint convex cells and solves a MAPF instance in each cell. Partitioning has two benefits: 1) fewer robots for each subgraph, and 2) decomposed instances can be solved in parallel. While the user can adopt any generic workspace partitioning approach, we propose a method that generates $Q$ convex polytopes. First, we use graph partitioning (KaHyPar [20]) to group the roadmap into $Q$ balanced (similar vertex numbers) subgraphs $\mathcal{G}_{m}\!\!=\!\!\left({V_{m},E_{m}}\right)$, for $m\!\!=\!\!1,\ldots,Q$. Balanced subgraphs lead to cells of similar volume and an even spread of robots.
We further enforce each cell, containing a subgraph, to be a convex polytope. Cell convexity prevents robots from penetrating into neighboring cells before exiting their current cells. We use soft-margin support vector machines (SVM) [21] to compute a hyperplane $H_{ml}$ between vertices of $\mathcal{G}_{m}$ and $\mathcal{G}_{l}$, and reassign misclassified vertices. Here, the subscript $l$ denotes the index of the subgraphs with vertices connecting to vertices of $\mathcal{G}_{m}$. The resulting set of hyperplanes forms the $m$-th cell, denoted as $P_{m}$.
@@ V-C Local Goal Generation
For navigating out of a cell, we generate candidate goal states on the faces between adjacent cells. For each cell $P_{m}$, we uniformly sample random local goals on the hyperplane $H_{ml}$ and add them as shared vertices to both $\mathcal{G}_{m}$ and $\mathcal{G}_{l}$. Despite being shared vertices, local goals generated by partition $P_{m}$ have in-edges from $P_{m}$ and out-edges to $P_{l}$ to avoid collision during cell transit. Thus, this part of the graph is directed. To enable parallel computation, there must be no communication between cells. Thus, the cell roadmap $\mathcal{G}_{m}$ is modified such that the planned paths are collision-free when crossing cells without information exchange between cells. The following properties should be satisfied:
P1: Robots avoid collision when stationary at vertices (local-goal or non-local-goal vertices) of different cells (generalized vertex-vertex conflict across cells), i.e., $\forall i,j,m\!\!\neq\!\!l\!:\!v^{i}_{m}\!\!\notin\!\!\operatorname{conVV}(v^{j}_{l})$, where the superscript in $v^{i}_{m}$ refers to the vertex index and the subscript refers to the cell index.
P2: Robots avoid collision when traversing edges between different cells (generalized edge-edge conflict across cells), i.e., $\forall i,k,m\!\!\neq\!\!l:e^{i}_{m}\!\!:=\!\!(v^{i}_{m},v^{j}_{m})\notin\operatorname{conEE}(e^{k}_{l})$, here the superscript in $e^{i}_{m}$ refers to the edge index.
@@
We depict violations of P1-P3 in Fig. 3b-3d. To prevent conflicts between stationary robots at local-goal and non-local-goal vertices across cells, we buffer the separating hyperplane by the inter-robot collision configuration, $\mathcal{C}_{col}$ (c.f. Fig. 3a), which is computed as $\mathcal{C}_{col}(\mathbf{p})=\mathcal{R}_{\mathcal{R}}(\mathbf{p})\oplus\mathcal{R}_{\mathcal{R}}(\mathbf{0})$, where $\oplus$ is the Minkowski sum. The buffering is achieved by modifying the offset $\mathcal{H}_{a}^{{}^{\prime}}=\mathcal{H}_{a}+\mathrm{max}_{\mathbf{y}\in\mathcal{C}_{col}(\mathbf{0})}\mathcal{H}_{\mathbf{n}}\!\cdot\!\mathbf{y}$ of the hyperplane, where $\mathcal{H}_{a}$ and $\mathcal{H}_{\mathbf{n}}$ are the offset and the normal vector of the hyperplane. Given the buffered hyperplanes, all non-local-goal vertices within the buffered region are removed, preventing collisions between stationary robots at local-goal and non-local-goal vertices. The local goals sampled on the hyperplane are confined within the blue region in Fig. 3a to avoid vertex-vertex conflicts between stationary robots at local goals of different separating hyperplanes. The buffering satisfies P1 between local-goal and non-local-goal vertices across cells and P1, P2, and P3 between non-local-goal vertices across cells. To satisfy P1 between local-goal vertices across cells, we uniformly randomly sample points on the hyperplane and reject those that violate P1. For P2 and P3 between local-goal and non-local-goal vertices, we add the connection between the sampled local-goal vertex to valid (no collision with the environment) non-local-goal vertices within a radius on both sides of the hyperplane. The local goal is removed if any connected edge violates P2 or P3 (see Fig. 3c, 3d). Otherwise, the edges are added to the cell roadmaps. Figure 4 depicts an exemplar partition.